OpenStudy (tylerbanks2001):

What is the solution to –4|–2x + 6| = –24?

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. −4(|−2x+6|)=−24 Step 1: Divide both sides by -4. −4(|−2x+6|) −4 = −24 −4 |−2x+6|=6 Step 2: Solve Absolute Value. |−2x+6|=6 We know either−2x+6=6or−2x+6=−6 −2x+6=6(Possibility 1) −2x+6−6=6−6(Subtract 6 from both sides) −2x=0 −2x −2 = 0 −2 (Divide both sides by -2) x=0 −2x+6=−6(Possibility 2) −2x+6−6=−6−6(Subtract 6 from both sides) −2x=−12 −2x −2 = −12 −2 (Divide both sides by -2) x=6 Answer: x=0 or x=6